OpenStudy (jenniferjuice):

Pleaseee help????😭 State the conditions on the values a and k so that the parabola y=a(x-h)^2+k opens downward and intersects the x-axis in two points

OpenStudy (anonymous):

a must be negative so that the parabola slopes downward and k must be greater than 0 so that the parabola intersects the x axis. \[a<0 & k>0\]

OpenStudy (michele_laino):

please, note that,the condition on a, is: \[a<0\] as @Jaguar1998 well wrote, so if I set: \[a=-\alpha ^{2}\] I can re-write your parabola as follows: \[y=-\alpha ^{2}(x-h)^{2}+k\] and it intersects the x-axis, if and only if: \[\alpha ^{2}(x-h)^{2}=k\] from which: \[x=h \pm \frac{ \sqrt{k} }{ \alpha } \] and as you can see, the solution exist if \[k \ge 0\]

OpenStudy (michele_laino):

oops...the solutions can exist if...

OpenStudy (anonymous):

Michel your answer is flawed in many ways. for one, the negative next to alpha squared must also be squared. Also if k is equal to zero then the graph only intersects the x axis at one point.

OpenStudy (michele_laino):

for your first question, I think you are wrong, because I have to solve this equation: \[-\alpha ^{2}(x-h)^{2}+k=0\] for your second question you are wrong, because the text ofquestion of @jenniferjuice doesn't specify if the points have to be distinct, so if k=0, we have to solve this equation:

OpenStudy (jenniferjuice):

sorry my keyboard was being stupid im on my laptop now

OpenStudy (michele_laino):

sorry, for your second question you are right! k>0 only @Jaguar1998

OpenStudy (jenniferjuice):

okay so a and k are 0?

OpenStudy (anonymous):

Michele, the author specified that the graph must intersect the x axis at two points. Also -alpha^2 does not equal a because you are changing the signs. (-alpha)^2 can equal a. but the negative sign must be in the parenthesis

OpenStudy (michele_laino):

a<0 and k>0

OpenStudy (anonymous):

my original answer :)

OpenStudy (michele_laino):

yes! youy are right! @Jaguar1998 if k=0, our parabola degenerates in a couple of lines!

OpenStudy (michele_laino):

please I wrote \[a=-\alpha ^{2}\] in order to work with a positive quantity @Jaguar1998

OpenStudy (michele_laino):

...or in order to explicit the negative sign